A new grill has a mass of 30.0 kg. You put 2.0 kg of charcoal in the grill. You burn all the charcoal and the grill has a mass of 30.0 kg. What is the mass in kg of the gases given off? (assume that the charcoal is pure carbon solid and that it burns completely in oxygen).
1
Expert's answer
2010-09-29T02:41:20-0400
Mass of carbon burnt = 2Kg = 2000g
Moles of Carbon burnt = 2000/12
=> Moles of CO2 produced = 2000/12
=> Mass of CO2 produced = 2000/12*44 = 2 333.33 g = 2.33 Kg